An individual who has automobile insurance from a certain company is randomly selected. Let Y be the number of moving violations for which the individual was cited during the last 3 years. The pmf of Y is the following.
y 0 1 2 3
p(y) 0.50 0.25 0.20 0.05
(a) Compute E(Y).
E(Y) =

(b) Suppose an individual with Y violations incurs a surcharge of $110Y2. Calculate the expected amount of the surcharge.
$
Answer #1
Concepts and reason

Mathematical expectation is the product of the probability of an event occurring and the value corresponding with the actual observed occurrence of the event.

Fundamentals

The expected value of the random variable can be defined as,

Ε(Y) =Σp(y)

And

Ε(Y?)=Σ» p(y)

EaY)= aE (Y)

(a)

The probability distribution table is as shown below:

1 y
| P(y)
0 11 12 13
0.5 0.25 0.2 0.05

Find the value of E(Y)

Ε(Υ)=ΣΥΡ(ν)
= (0 x 0.5) + (1x 0.25) + (2 x 0.20) + (3 x 0.05)
= 0 + 0.25 + 0.40 + 0.15
= 0.80

(b)

Suppose an individual with violations incurs a surcharge of $110 , the expected amount of the surcharge is given by,

E($11072) = 110 E(Y)
=110 y’p(y)
=110{(0’ 0.5) +(1* 0.25) + (22 x 0.20) + (32 < 0.05)
=110{0 +0.25 +0.80 +0.45}
= $165

Ans: Part a

The expected value of the number of moving violations for which the individual was cited during the last 3 years is

Part b

The expected amount of surcharges is $165.